# AI-PACE: A Framework for Integrating AI into Medical Education

### v1

@@ 5 Implementation Strategy
Figure 2: A proposed approach to implementing an AI curriculum into an existing program.
@@ 6 Discussion
The rapid evolution of AI in medicine has outpaced the development of comprehensive educational frameworks, creating a "preparedness gap" that threatens the safe integration of these technologies into clinical practice. The AI-PACE framework addresses these limitations by offering a unified, spiral curriculum model that evolves with the learner.
6.1 The "Integration vs. Addition" Debate
A primary barrier to AI education is the "crowded curriculum" problem. Medical school and residency schedules are already saturated (Wartman, 2019). Integrating AI concepts into existing courses rather than creating new courses can help address this challenge (Grunhut et al., 2022). For example, the cognitive distinction between deterministic algorithms and probabilistic machine learning should be taught alongside EBM during pre-clerkship years. Similarly, the psychomotor skill of "human-in-the-loop" verification should be practiced during existing clinical clerkships, explicitly avoiding the “fragmented bootcamp” model that detaches technology from the daily clinical workflow. This integration model mirrors how EBM itself was once a novel addition that is now woven into the fabric of clinical reasoning.
@@ 6.2 The Resource Challenge: Faculty and Funding
Implementing AI-PACE requires overcoming significant systemic hurdles. Many medical educators lack expertise in AI and data science (Grunhut et al., 2022; Chan and Zary, 2019). This can be addressed through faculty development programs, collaboration with computer science departments, and recruitment of clinician-informaticians. Briganti and Le Moine highlight that many schools are addressing this by associating medical curricula with engineering curricula (Briganti and Le Moine, 2020). Furthermore, the fast pace of AI development risks making curricular content quickly outdated (Charow and others, 2021). Institutions must leverage interdisciplinary partnerships, breaking down the silos between Computer Science, Informatics, and Medicine, to share resources and ensure content remains current (Masters, 2020).
@@ 6.3 Validation and Future Directions
As medical education adopts AI curricula, we must rigorously validate their impact. Future research should move beyond self-reported confidence (Kirkpatrick Level 1) to measure changes in behavior (Level 3) and patient outcomes (Level 4) (Kirkpatrick and Kirkpatrick, 2006; Yardley and Dornan, 2012). Does an AI-literate resident make safer decisions when using a diagnostic support tool? By standardizing the educational model through AI-PACE, we can begin to generate the multi-institutional data needed to answer these questions. This is critical because we are learning that relying on GenAI for efficiency may come at a cost of skill formation (Shen and Tamkin, 2026).
@@ 7 Conclusion
The integration of AI into healthcare necessitates corresponding changes in medical education. Our analysis of the literature reveals a significant gap between the rapid adoption of AI in clinical practice and the preparation of physicians to use these technologies effectively. While a limited number of educational initiatives exist, comprehensive curriculum frameworks are largely absent or fragmented. We have identified key competencies across cognitive, psychomotor, and affective domains that should guide AI education for medical professionals. Effective curricula should balance technical knowledge with clinical applications and ethical considerations, be integrated longitudinally throughout medical training, and involve interdisciplinary collaboration.
The AI-PACE framework presented in this paper offers a starting point for medical educators seeking to develop AI curricula. Future research should focus on implementing and evaluating such curricula, developing specialty-specific educational approaches, and creating faculty development programs to support AI education. As AI continues to transform healthcare, ensuring that physicians are adequately prepared to leverage these technologies while maintaining humanistic, patient-centered care remains a critical challenge for medical education.

### v2

@@ 5 Implementation Strategy
Figure 2: A six-step implementation cycle for AI curriculum development. Steps progress from defining purpose, establishing objectives, designing/organizing content, teaching strategies, and evaluation. The evaluation step is continuous rather than terminal, reflecting the accelerated pace at which AI capabilities and clinical applications evolve.
@@ 6 Discussion
The rapid evolution of AI in medicine has outpaced the development of comprehensive educational frameworks, creating a “preparedness gap” that threatens the safe integration of these technologies into clinical practice. The AI-PACE framework is designed to address these limitations through a spiral curriculum model that evolves with the learner.
6.1 The “Integration vs. Addition” Debate
A primary barrier to AI education is the “crowded curriculum” problem. Medical school and residency schedules are already saturated (Wartman, 2019). Integrating AI concepts into existing courses rather than creating new courses can help address this challenge (Grunhut et al., 2022). For example, the cognitive distinction between deterministic algorithms and probabilistic machine learning should be taught alongside EBM during pre-clerkship years. Similarly, the psychomotor skill of “human-in-the-loop” verification should be practiced during existing clinical clerkships, explicitly avoiding the “fragmented bootcamp” model that detaches technology from the daily clinical workflow. This integration model mirrors how EBM itself was once a novel addition that is now woven into the fabric of clinical reasoning.
At the University of California, Davis School of Medicine, Health System Science was introduced as a new longitudinal pillar through the I-Explore curriculum reform folded into existing courses rather than added as standalone credit hours (UC Davis School of Medicine, 2021). AI literacy could follow the same path, with cognitive foundations embedded in Biostatistics and EBM coursework, psychomotor skills practiced during existing clinical clerkships, and affective competencies addressed through the documentation and supervision structures already present in clinical training. At the GME level, Preiksaitis demonstrates that supervising resident AI use can be embedded into existing case presentations and clinical rounds through structured prompts, without requiring dedicated curriculum time (Preiksaitis, 2026).
@@ 6.2 The Resource Challenge: Faculty and Funding
Implementing AI-PACE requires three categories of institutional investment. First, faculty development: many medical educators lack expertise in AI and data science (Grunhut et al., 2022; Chan and Zary, 2019), which can be addressed through structured upskilling pathways, recruitment of clinician-informaticians, and partnerships with engineering curricula (Briganti and Le Moine, 2020). Second, interdisciplinary infrastructure: breaking down silos between Computer Science, Informatics, and Medicine is necessary to keep content current as AI capabilities evolve rapidly (Charow and others, 2021; Masters, 2020); institutions with existing health informatics graduate programs may already have this partially in place. Third, technical access: students frequently lack the same licensed access to AI tools available to clinical staff, as licensing costs and IT compliance support are typically budgeted for employees rather than trainees. Proactive coordination between medical school administration and health system IT governance is required to close that gap.
@@ 6.3 Validation and Future Directions
As medical education adopts AI curricula, we must rigorously validate their impact. Future research should move beyond self-reported confidence (Kirkpatrick Level 1) to measure changes in behavior (Level 3) and patient outcomes (Level 4) (Kirkpatrick and Kirkpatrick, 2006; Yardley and Dornan, 2012). Does an AI-literate resident make safer decisions when using a diagnostic support tool? By standardizing the educational model through AI-PACE, we can begin to generate the multi-institutional data needed to answer questions like this. This work is also timely given emerging evidence that relying on GenAI for efficiency may come at a cost of skill formation (Shen and Tamkin, 2026). This is particularly relevant given emerging concerns about deskilling, mis-skilling, and never-skilling as physicians increasingly offload cognitive tasks to AI tools (Natali et al., 2025; Berzin and Topol, 2025).
To move from framework to evidence, we are currently piloting AI-PACE through a 4-week elective course for medical students at the University of California, Davis. Pre- and post-intervention AI preparedness will be measured using the Medical Artificial Intelligence Readiness Scale for Medical Students (MAIRS-MS), a validated instrument developed for this purpose (Karaca et al., 2021). Results from this pilot are intended as a follow-up case study to this paper. Beyond the US context, international adaptation of AI-PACE represents an important next step, one that would likely require a broader consensus methodology, such as a Delphi study, to account for the variation in training structures and accreditation requirements across health systems.
Neither the Liaison Committee on Medical Education (LCME) nor the Accreditation Council for Graduate Medical Education (ACGME) currently have binding AI-specific standards, but the field is moving in that direction. A 2025 Josiah Macy Jr. Foundation conference co-sponsored by AAMC and ACGME explicitly recommended implementing AI curricula and competencies and modifying accreditation processes to account for AI (Josiah Macy Jr. Foundation, 2025). AI-PACE is designed to map onto this trajectory, providing institutions with a structured framework for responding to emerging accreditation expectations rather than waiting for prescriptive requirements that do not yet exist.
6.4 Limitations
Several limitations of this work should be acknowledged. First, article screening was conducted by a single reviewer (S.P.M.), which introduces the potential for selection bias. A formal systematic review with blinded, multi-reviewer screening was outside the scope of this framework development effort. Second, the review was intentionally restricted to medical training programs in the United States, reflecting the immediate practical need that motivated this work: developing a pilot course for US medical students. This US-centric focus limits the generalizability of AI-PACE to international contexts, where training structures, accreditation requirements, and pedagogical approaches can vary considerably. There is the potential for international adaptation of the framework, which could be achieved through a consensus methodology such as a Delphi study. Third, this manuscript cites formal technical reports and model cards from frontier AI laboratories alongside peer-reviewed literature when characterizing current model capabilities. This reflects a structural limitation of the field: the 9–12-month peer-review publication cycle means that empirical evaluations of LLM performance often assess models one or two generations behind those currently in clinical use. System cards and model cards, while not peer-reviewed, represent the timeliest available documentation of capability boundaries and safety evaluations for recently released models. We acknowledge this as a methodological constraint and encourage readers to weigh these sources accordingly.
@@ 7 Conclusion
The integration of AI into healthcare necessitates corresponding changes in medical education. Our analysis of the literature reveals a significant gap between the rapid adoption of AI in clinical practice and the preparation of physicians to use these technologies effectively. While a limited number of educational initiatives exist, comprehensive curriculum frameworks are largely absent or fragmented. We have identified competencies across cognitive, psychomotor, and affective domains that should guide AI education for medical professionals. Effective curricula should balance technical knowledge with clinical applications and ethical considerations, be integrated longitudinally throughout medical training, and involve interdisciplinary collaboration.
The AI-PACE framework presented in this paper offers a starting point for medical educators seeking to develop AI curricula. Future research should focus on implementing and evaluating such curricula, developing specialty-specific educational approaches, and creating faculty development programs to support AI education. As AI continues to transform healthcare, ensuring that physicians are adequately prepared to leverage these technologies while maintaining humanistic, patient-centered care is among the most pressing challenges facing medical education.
Acknowledgements
This work was supported by a grant provided to CITRIS Health at UC Berkeley from The First Affiliated Hospital, Sun Yat-Sen University (grant number A25-2059-002), and by the National Center for Advancing Translational Sciences, National Institutes of Health (grant number UL1 TR001860).
S.P.M. thanks Melanie L. McGrath, PhD, LAT, ATC, for her assistance with developing the manuscript in its earliest form and for her steadfast support.
Author Contributions
S.P.M. conceptualized the framework, wrote the original draft, and designed the figures. K.K.K., K.J., H.W., and N.A. critically reviewed and edited the manuscript. All authors have read and approved the final manuscript.
Competing Interests
The authors declare no competing interests.